# Is fixed-node diffusion quantum Monte Carlo reproducible?

### v1

@@ 1 Introduction
There exist several codes that have been under active development for decades, each of which makes somewhat different algorithmic and implementation choices, such as different Jastrow factors. This diversity raises open questions on the reproducibility of FN-DMC. If FN-DMC is to be widely accepted as a highly accurate reference method, it is important that consistent results can be obtained from these different FN-DMC codes. With this goal in mind, the present letter represents a collaborative effort among the users and developers of eleven distinct FN-DMC codes, to rigorously assess the reproducibility of FN-DMC.
A key obstacle to the reproducibility of FN-DMC comes from the use of non-local pseudopotentials (NLPPs), which increase the efficiency of the method for systems with heavy atoms. [95, 96, 97] NLPPs allow one to solve the Schrödinger equation solely for the valence electrons, by substituting the full local nuclear potential with a smooth non-local potential near the nuclei. In general, NLPPs hinder reproducibility in electronic structure methods, as NLPPs constructed in different ways can lead to somewhat different predictions. This issue is even more challenging in FN-DMC because non-local pseudopotential operators create an additional sign problem in the projector beyond the one that is always present for fermionic calculations. To avoid this sign problem, these operators must be “localized”,[98] or at least partially localized,[99] on a wave-function. A natural choice is to localize them on the trial wave-function , introducing a dependence on both the determinantal and the Jastrow components of the wave function. Since the Jastrow factor is different in the different codes and its parameters are stochastically optimized, yielding possible noise and reproducibility issues, some authors choose to localize only on the determinantal component. [100, 96, 101, 102, 103] This removes the dependence on the Jastrow factor at the cost of losing the desirable property that the treatment of the pseudopotential is exact in the limit of exact . To summarize, there are currently four localization schemes: the locality approximation (LA)[104, 98], the T-move (TM) approximation[99, 105, 106], the determinant locality approximation (DLA)[100, 96, 101, 102, 103], and the determinant T-move (DTM) approximation[103]. These four schemes (LA, TM, DLA and DTM) result in somewhat different projected wave-functions and therefore different total energies of physical systems.
@@ 2 Results and Discussion
where is the number of codes for which the localization method is evaluated. The mean and the variance of the energy for the distribution are respectively:
@@
The probability distributions are plotted in Fig. 2. When the LA is employed, the probability distribution is spread across a large energy range of , with a standard deviation of meV. The agreement across different codes significantly improves with the TM, DLA and DTM schemes, with the probability distributions showing a quite localized peak (standard deviation of ca. meV or less) centered on meV, meV and meV respectively. The DTM scheme gives the narrowest distribution, centered on meV, with a standard deviation of ca. 1 meV, but since only four out of the eleven codes implemented DTM this is not of much significance. Overall, the analysis of the probability distributions showcases that algorithms more sophisticated than LA need to be employed to guarantee reproducibility among different FN-DMC codes.
@@
To conclude, this work establishes FN-DMC as a robust reference method by demonstrating reproducibility across 11 popular codes with varying implementations. This work marks a significant milestone for FN-DMC and the greater electronic structure theory community, giving confidence in the multidisciplinary array of applications of FN-DMC. Crucially, we show that reliable and reproducible FN-DMC predictions can be achieved when using the TM, DLA, or DTM approximation schemes, rather than the LA scheme, for pseudopotential simulations while ensuring convergence with respect to the time step. Notably, we are able to achieve agreement to within a standard deviation of meV on the interaction energy of the methane-water complex — which is approximately two hundred thousand times smaller than the total energy of the dimer. This achievement lays the groundwork for more extensive studies in the future, including calculations on the physical and chemical properties in the condensed phase.

### v2

@@ 1 Introduction
There are numerous QMC codes currently used for research, many of which have been under development for over a decade. Each makes somewhat different algorithmic and implementation choices, such as use of different Jastrow factors and methods for evaluating single particle orbitals. As detailed below, we compare eleven of them in this study. This diversity raises open questions on the reproducibility of FN-DMC. If FN-DMC is to be widely accepted as a highly accurate reference method, it is important that consistent results can be obtained from these different FN-DMC codes. With this goal in mind, the present letter represents a collaborative effort among the users and developers of eleven distinct FN-DMC codes, to rigorously assess the reproducibility of FN-DMC.
A key obstacle to the reproducibility of FN-DMC comes from the use of non-local pseudopotentials (NLPPs), which increase the efficiency of the method for systems with heavy atoms. [95, 96, 97] NLPPs allow one to solve the Schrödinger equation solely for the valence electrons, by substituting the full local nuclear potential with a smooth non-local potential near the nuclei. In general, NLPPs hinder reproducibility in electronic structure methods, as NLPPs constructed in different ways can lead to somewhat different predictions. NLPPs are a potential source of non-reproducibility in FN-DMC even when the same NLPPs are used, because non-local pseudopotential operators create an additional sign problem in the projector beyond the one that is always present for fermionic calculations. To avoid this sign problem, these operators must be “localized”,[98] or at least partially localized,[99] on a wave-function. A natural choice is to localize them on the trial wave-function , introducing a dependence on both the determinantal and the Jastrow components of the wave function. Since the Jastrow factor is different in the different codes and its parameters are stochastically optimized, yielding possible noise and reproducibility issues, some authors choose to localize only on the determinantal component. [100, 96, 101, 102, 103] This removes the dependence on the Jastrow factor at the cost of losing the desirable property that the treatment of the pseudopotential is exact in the limit of exact . To summarize, there are currently four localization schemes: the locality approximation (LA)[104, 98], the T-move (TM) approximation[99, 105, 106], the determinant locality approximation (DLA)[100, 96, 101, 102, 103], and the determinant T-move (DTM) approximation[103]. These four schemes (LA, TM, DLA and DTM) result in somewhat different projected wave-functions and therefore different total energies of physical systems.
@@ 2 Results and Discussion
where is the number of codes for which the localization method is evaluated. The mean, , and the variance, , of the energy for the distribution are respectively:
@@
The probability distributions are plotted in Fig. 2. When the LA is employed, the probability distribution is spread across a large energy range of , with a standard deviation of meV. The agreement across different codes significantly improves with the TM, DLA and DTM schemes, with the probability distributions showing a quite localized peak (standard deviation of ca. meV or less) centered on meV, meV and meV respectively. The DTM scheme gives the narrowest distribution, centered on meV, with a standard deviation of ca. 1 meV, but since only four out of the eleven codes implemented DTM this is of limited significance. Overall, the analysis of the probability distributions showcases that algorithms more sophisticated than LA need to be employed to guarantee reproducibility among different FN-DMC codes.
@@
Finally, we focus on the FN-DMC total energies of the methane-water dimer and its constituent monomers, which are the fundamental quantities entering the computation of the interaction energy. In Fig. 4, we report the probability distribution of the total energies extrapolated to zero-time step. As in the case of the interaction energy, we find that the total energies computed in the TM, DLA, and DTM approximations differ much less among the codes than when the LA is employed. Their distributions are significantly narrower, displaying standard deviations in a range from 2.5 to 10 times smaller than the LA case (e.g., in the water molecule , and in the methane monomer ). Moreover, the standard deviations s of the TM, DLA and DTM total energy distributions are close to the theoretical minimum allowed by the precision of the performed FN-DMC simulations, as s are mostly determined by the stochastic error associated to the FN-DMC energy evaluations (between and Hartree, see SI), so the first term on the right hand side of Eq. 4. This behavior is expected for the DLA and DTM schemes that depend only on the determinant part of the wave functions (identical in all calculations). Remarkably, despite using different Jastrow factors, all codes yield very similar extrapolated total energies even with the TM scheme, which has the desirable property of treating the pseudopotential exactly in the limit of an exact .
@@ 3 Methods
4 Summary and Conclusions
In this work, we investigated the reproducibility of FN-DMC calculations across 11 popular QMC codes which differ in the details of the algorithms used. This study represents a significant collaborative effort, involving more than 300 FN-DMC calculations, spanning 11 codes, multiple DMC time steps, and different pseudopotential localization schemes. Our results establish FN-DMC as a robust reference method by demonstrating its reproducibility.
In particular, we conducted a thorough analysis of two key obstacles to FN-DMC reproducibility, namely the use of NLPPs and finite time-step bias. We systematically compared four localization schemes, LA, TM, DLA, and DTM, for the interaction energy of the methane-water dimer and found that agreement across all eleven codes is achieved in the limit of zero time step when employing the TM, DLA, and DTM approximations. Notably, we achieve agreement within a standard deviation of meV on the interaction energy of the methane-water complex, approximately two hundred thousand times smaller than the total energy of the dimer. Larger discrepancies are observed with the LA scheme. Moreover, we show that even the total energies with the TM, DLA, and DTM schemes have a standard deviation among the codes which is smaller than meV. This agreement further reinforces the reproducibility of FN-DMC.